Derive the value of flux density at the specified point along the axis due to circular coil carrying current.


1
Expert's answer
2022-01-25T14:07:47-0500

"B=\int_0^{2\pi a}dB\sin\theta=\frac {\mu_0 }2\frac{Ia^2}{(x^2+a^2)^{\frac 32}}."